OpenStudy (anonymous):

A cube has side length measured to be 12 cm. Find the (linearly) approximate error in volume if the error in side length measurement is Δs.

OpenStudy (anonymous):

\[ V=r^3 \]So \[ dV = \frac{dV}{dr} dr=3r^2dr \]

OpenStudy (anonymous):

Change those \(r\)s into \(s\)s...

OpenStudy (anonymous):

So we know \(s=12cm\).\[ dV=3(12)^2dr = 3(144)dr=432dr \]

OpenStudy (anonymous):

Thanks! Is that the final answer?

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

Mm, what am I supposed to do after?

OpenStudy (anonymous):

You only have how 'off' it is.

OpenStudy (anonymous):

The problem is that I'm not quite sure what kind of error they're looking for.

OpenStudy (anonymous):

If they want relative error, then we want\[ dV/V = 432\Delta s/(12)^3 \]

OpenStudy (anonymous):

Ahh, okay, thanks.

OpenStudy (anonymous):

That gets us\[ \frac{\Delta V}{V} \approx \frac{dV}{V} =\frac{3\Delta s}{12}=\frac{\Delta s}{4} \]

OpenStudy (anonymous):

Okay thanks! :)

OpenStudy (anonymous):

got a way of checking the answer?

OpenStudy (anonymous):

Nope, but I'm sure it's right.